Plan for September 2018 Startup Observing


Times listed are times on-sky without any weather contingency.

Decision Tree

Is Instrument Mounted? 
   N: Check next instrument
   Y: Is Daytime Checkout Completed?
      N: Check next instrument
      Y: Is On-sky Wake-Up Completed?ovms_hdfdirect.py
         N: Do On-sky Wake-Up
         Y: Is it photometric?
            N: Are Commissioning Tests Completed?
               N: Do Commissioning tests.
               Y: Do other software or engineering tests.
            Y: Is On-sky Calibration Completed?
               N: Do calibration observations.
               Y: Is Science Verification Completed?
                  N: Do verfication observations.

Followup tasks for E-time and D-time in second half of September

Note that these are all things that we want to do, but as of Friday afternoon weren't quite ready. None of them need significant time on sky, but may need significant work off-sky.
  • PCS patch for IT 7380 (waiting for Petr to discover the bug in the co-pointing initialization)
  • Apply binodal correction to SX Gregorian (waiting for Doug to analyze the data)
  • Apply updated pointing model (waiting for Dave to process the LUCI pointing data)
  • Apply AGw4 transform (waiting for Doug, who was having problems with the analysis last week)
  • Apply AGw2 transform (waiting for Doug)(less urgent as AGW2 already has a decent transform)

And then some things that could use sky time now:
  • collect pointing data with MODS/MODS (~2.5 hours, any seeing)
  • check MODS1 guide probe wrt/rotator center (0.5 hour in decent seeing)
  • collect MODS1 transform data (Olga, 2 hours in good seeing)
  • refine AGw4 lenslet alignment (John, in parallel to other LUCI activities)
  • SX SOUL on-sky?
  • Test sending LBC + LUCI presets via scripts (Olga, a bad-seeing, bad-weather task)

And then some other things that we'll need to work on later (beyond September):
  • DX M3 alignment at other focal stations
    • for PFU (high)
    • for L-N (high)
    • for LBTI (high)
  • Temperature-free collimation model (from Yang's analysis)
  • More pointing data after the initial models are tested
  • Test new version of GCS

General Scope of Activities

  • Telescope wake-up
  • Instrument wake-up (LBCs, MODS and LUCIs)
  • TCS Software Build 2018B
  • M3 alignment
  • AGW alignment
  • AO on-sky tests of SOUL (FLAO and LBTI)
  • Instrument Calibrations (LBCs, MODS, and LUCIs)
  • Training of new telescope operator

List of On-sky Tasks

  • DONE Telescope wake-up
  • LBC wake-up checks (high)
    • DONE Collimate WRS/DOFPIA - high (4 x 0.25 hrs)
    • DONE Superfocus sequence -PA=0,180 high (1 hr)
    • DONE Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr)
    • Test Science OB - guiding - medium (1 hr)
      • DONE Sidereal
      • DONE Non-Sidereal - IT 7387 guiding with NS not working. Otherwise NS commands successfully sent.
    • DONE Standard photometric fields (check centering of pre-made OBs) - medium (1 hr)
    • DONE Spectro-photometric standards
    • Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete)
      • B-Bessel (PA=0,180)
      • z-sloan (PA=0,180)
      • Uspec (PA=0,180)
      • I-Bessel (PA=0,180)
      • g-sloan (PA=0,180) - need to repeat (3/5 OK PA=0, 5/5 OK PA=180)
      • i-sloan (PA=0,180)
      • V-Bessel (LBC-B, LBC-R) - (PA=0,180)
      • r-sloan (LBC-B, LBC-R) - (PA=0,180)
      • U-Bessel (PA=0,180)
      • R-Bessel (PA=0,180)
    • TMS pointing model (DJT email 20180208) - log
  • MODS on-sky checkout (high) - (MODS 2 installed Sep 5, MODS 1 installed Sep 6) - MODS2 functional checkouts start Friday Sep 7
    • DONE Rotator Center measure (can use same field as for astrometric masks - StoneO or StoneA) - high (0.5hr)
    • DONE Measure Rotator zero point (mask alignment, masks for StoneO and StoneA) - high (0.5 hr)
    • AGW Transform data - ~1" seeing - high (2hours to collect data on-sky, time during day to analyze, ~0.5 hr-1hr to verify on-sky)
    • DONE Spectrophotometric standard all modes - medium
      • Feige 110 Dual Grating MODS-1, MODS-2 Blue airmass 1.7 and 1.3; G191B2b MODS-1 Dual airmass 1.12
    • non-sidereal guiding (imaging and longslit) - low (0.5 hrs)
    • pointing model data collection - medium/high (1-2 hrs)
    • Test Science OB - DONE longslit (MODS1 and MODS2) and DONE imaging (MODS-1 only) low - (1hr)
  • M3 alignment
    • DONE for LFBG/RFBG (high)
    • for PFU (high)
    • for L-N (high)
    • for LBTI (high)
  • AGW4 checkout and alignment at LFBG (critical) (started 20180903 UT)
    • DONE initial wake-up
    • DONE general wfs alignment - nightly work
    • DONE align hotspot to rotator center
    • DONE adjust M3 to minimize pupil wobble
    • understanding the remaining bugs and features
  • AGW2 checkout and alignment check at RFBG (high) (started 20180903 UT)
    • DONE initial wake-up
    • DONE general wfs alignment
    • DONE align hotspot to rotator center
    • DONE adjust M3 to minimize pupil wobble
  • DONE measure LFBG/RFBG field aberrations (high)
  • DONETCS Build 2018B on-sky verification (medium) - 20180905 UT
  • PEPSIPFU wake-up checks (medium)
  • bent Gregorian TMS tests (low) - postponed by TMS late parts delivery
  • LUCI-1 and LUCI-2
    • DONE Rotation Zeropoint (Astrometric masks - to be done after Sep 7-LUCI-1 and Sep-10-LUCI-2) - high (1hr)
    • DONE Check Orientation on-sky - high (1hr)
    • DONE Image Pointing and Acquisition Check - high (1 hr)
    • DONE AGW-4 and DONE AGW-2 Transform - high (2x2hrs - requires initial data, then followup to confirm transform)
    • DONE Longslit and DONE Imaging check (dither along slit, offsets and dithering image tests) - high (1 hr)
    • DONE non-sidereal check imaging - high (0.5hr)
    • non-sidereal check longslit - low (0.5 hr)
    • Twilight flats zJHKKs (narrow band if time) - low (2 hrs, twilght 1hr each)
    • Test Science OB -
      • DONE imaging - medium (1 hr)
      • DONE longslit - medium (0.5 hr)
    • LUCI-2 DX AO test images - high (0.5 hr)
    • LUCI-2 DX ESM test images - medium (0.5 hr - spectroscopy highest priority, H and J imaging 2nd priority)
    • DONE pointing model data collection - medium/high (1-2 hrs)

List of Rainy Night Tasks

  • DONE AGW-4 Troubleshooting (https://lbt.issuetrak.com/CSIssue_View.asp?IssueNbr=7377)
  • AO calibration
  • LBC Checks
    • DONE Filter and Rotator Functional tests (filter test program did not move filter wheel)
  • LUCI calibration
    • DONE General Functional Checkout
    • DONE MOS functional Checkout
    • DONE Internal Focus Check
    • DONE Field centering and pupil alignment
      • LUCI-2/DX re-checked once M3 alignment completed
    • Minimum flexure calibration
    • or complete flexure calibration
    • DONE Grating tilt calibration
    • Imaging flats, spectroscopic cals (in standard/common settings) - low
  • Test ARGOS Cal Source for AGw alignment purposes
  • Test updated LBTtools package (JMH hasn't done it yet)
  • bent Gregorian TMS tests (after TMS is reinstalled)
  • MODS closed-dome checkout

Nighttime Calendar Schedule (2 - 13 September 2018 UT)

Night 1 (20180902 UT)

Observers: J. Hill, B. Rothberg
  • LOST TO WEATHER

Night 2 (20180903 UT)

Observers: J. Hill, B. Rothberg
  • DONE Telescope wake up check
  • AGW4 on-sky alignment - Most of Night lost to weather

Night 3 (20180904 UT)

Observers: J. Hill, B. Rothberg
  • AGW4 checkout and alignment at LFBG (critical) [4-5 hours] - progress made, not complete
  • LBC wake-up check
    • DONE Collimate WRS/DOFPIA - high (4 x 0.25 hrs)
    • DONE Superfocus sequence - high (1 hr)
    • DONE Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr)
    • Test Science OB - guiding - medium (1 hr)
      • DONE sidereal targets
      • DONE non-sidereal targets - troubleshooting guiding + NS w/o ephemeris

Night 4 (20180905 UT)

Observers: J. Hill, B. Rothberg
  • DONE TCS Software Build 2018B with P. Kubanek
  • AGW4 collimation work (not complete, but progress made)

Night 5 (20180906 UT)

Observers: J. Hill, B. Rothberg
  • AGW4 collimation work - progress
  • DONE AGW4 and AGW2 probe alignment
  • LBC Tasks Remaining:
    • LBC Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete) - if clear before sunset - configure for LBCs
      • DONE B, z PA=0, 180
  • LUCI Tasks-On-Sky - Preliminary to test dithering, guiding (DX only - SX collimation still a work in Progress, filter changes)
    • DONE Image Pointing and Acquisition Check - high (1 hr) - Functional Check
    • DONE Imaging check (dither along slit, offsets and dithering image tests) - high (1 hr) - Functional Check

Night 6 (20180907 UT)

Observers: J. Hill, B. Rothberg
  • AGW4 collimation work - In Progress
  • DONE SX and DX M3 alignment for pupil wobble (John and Barry)
  • DONE Field Aberration Data taken (DX Good, SX we are unsure - D. Miller analysis in progress)
  • LBC Tasks Remaining:
    • LBC Standard photometric fields (check centering of pre-made OBs) - medium (1 hr) - some progress,
      • DONE LBC Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete) - if clear before sunset - configure for LBCs
        • Uspec and I-Bessell at PA=0 and PA=180, i-sloan at PA=0, g-sloan (3/5) at PA=0
      • DONE Troubleshoot IT 7387 (guiding with NS tracking) - test will be to uncheck guiding in OB before rerunning observations from last night.
        • Data Taken

Night 7 (20180908 UT)

Observers: J. Hill, B. Rothberg
  • AGW4 collimation work - In Progress
  • DONE Field aberration data SX and DX
  • LUCI Tasks-On-Sky - DONE Requires M3 alignment for proper checks
    • AGW-4 and DONE AGW-2 Transform - high (2x2hrs - requires initial data, then followup to confirm transform)
    • DONE Check Orientation on-sky - high (1hr)
    • LBC Tasks remaining:
      • DONE Troubleshoot IT 7387 (guiding with NS tracking) - test will be to uncheck guiding in OB before rerunning observations from last night. - This time NS with guiding checked using Horizons Database

Night 8 (20180909 UT)

Observers: J. Hill, B. Rothberg
  • DONE AGW4 Collimation Work - In Progress
  • LUCI Tasks On-Sky
    • DONE AGW4 (and AGW2) Transform - high (2hrs)
    • DONE non-sidereal check (imaging) - High (0.5 hr)
    • Test Science OB Longslit - G210 N3.75, DONE N1.8; G200, N1.8 - high (1hr)
  • MODS Tasks (depends on cool-down progress)
    • DONE AGW 5/6 - closed dome check that works (powers on/light on CCD)
    • DONE AGW 5/6 - check on sky, lenslets aligned, collimation works, rotation center check
  • LBC - Continue Skyflats in Twilight

Night 9 (20180910 UT)

Observers: J. Hill, B. Rothberg, O. Kuhn
  • MODS Tasks On-Sky
    • DONE Rotator Center measure (can use same field as for astrometric masks - StoneO or StoneA) - high (0.5hr)
    • DONE Measure Rotator zero point (mask alignment, masks for StoneO and StoneA - high (0.5 hr)
    • DONE Spectrophotometric standard all modes - medium (priority for photometric sky)
    • DONE MODS-1 Imaging test (MODS-2 TBD)
    • DONE MODS test longslit OB

Night 10 (20180911 UT)

Observers: B. Rothberg, J. Power
  • LUCI Tasks On-Sky
    • DONE Zero point rotator check with new astrometric masks - high (0.5hr) - only tested on DX side
    • DONE MS/AFC tests - confirmation/checks of mask and software
  • LBCs
    • DONE Continue Skyflats in Twilight
    • DONE Photometric Standards
    • DONE Science OB - 3 OBs with collimation in between - exercising science-like
    • DONE Proper-motion co-pointing script checks (multiple tries)

Night 11 (20180912 UT)

Observers: B. Rothberg, J. Hill, and special guest S. Allanson
  • LUCi
    • DONE Zero Point rotator check with astrometric masks - need to collimate with IDL on SX
    • DONE Pointing Model observations - PtStars IRAF task (manual presets and manual image taking)
  • MODS-2 Blue
    • On-Sky tests - imaging and longslit
  • LBCs
    • Continue SpecPhot and Phot standards (if clear)
    • Science OB from UT 20180911 (if clear)
    • Run through filters (IC 342) if not clear
    • DONE Twilight Flats

Nighte 12 (20180913 UT)

Observers: B. Rothberg, J.Hill and special guest S. Allanson
  • MODS (1 or 2 or both)
    • AGW Transform
    • non-sidereal guiding (imaging and longslit) - low (0.5 hrs)
    • pointing model data collection - medium/high (1-2 hrs)
    • Test Science OBs (MODS-2 Blue if only one available)
  • LUCI
    • NS longslit acqu and data
    • AGW-4 work
    • AGW transforms

-- %USERSIG{JohnHill - 2018-07-11}%
Topic revision: r61 - 23 Oct 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback